### v1

# Unsupervised Multimodal Graph-based Model for Geo-social Analysis

Abstract
The systematic analysis of user-generated social media content, especially when enriched with geospatial context, plays a vital role in domains such as disaster management and public opinion monitoring. Although multimodal approaches have made significant progress, most existing models remain fragmented, processing each modality separately rather than integrating them into a unified end-to-end model. To address this, we propose an unsupervised, multimodal graph-based methodology that jointly embeds semantic and geographic information into a shared representation space. The proposed methodology comprises two architectural paradigms: a mono graph (MonoGrah) model that jointly encodes both modalities, and a multi graph (MultiGraph) model that separately models semantic and geographic relationships and subsequently integrates them through multi-head attention mechanisms. A composite loss, combining contrastive, coherence, and alignment objectives, guides the learning process to produce semantically coherent and spatially compact clusters. Experiments on four real-world disaster datasets demonstrate that our models consistently outperform existing baselines in topic quality, spatial coherence, and interpretability. Inherently domain-independent, the framework can be readily extended to diverse forms of multimodal data and a wide range of downstream analysis tasks.
keywords:
Machine learning, Artificial intelligence, Unsupervised machine learning, Natural language processing, Social media analysis, Disaster management
1 Introduction
Social media platforms continuously generate vast amounts of data, including posts and comments, likes, shares, images, and videos. This content is rich, diverse, and real-time, offering a valuable resource for understanding how people interact, express opinions, form communities, and respond to events. The widespread adoption of smartphones with built-in location-tracking capabilities has added another dimension to this landscape: geographic context. By enabling users to share their locations, these devices have facilitated the integration of geospatial data into online content, enhancing the contextual depth of social media analysis. A growing body of research underscores the importance of geo-location data in social media research. Prior studies have demonstrated its utility across a wide range of applications, from disaster response [goodchild2010crowdsourcing] to public health monitoring [Signorini2011]. In particular, geo-referenced content from micro-blogging platforms has been leveraged to model and monitor phenomena such as earthquakes [Havas2021Portability], floods [Huang2018Flood], refugee movements [Havas2021Refugee], and disease outbreaks [Stolerman2023Disease].
Previous social media research has often been focusing on semantic topic modeling, a fundamental analytical algorithmic technique to uncover latent thematic structures within large volumes of user-generated text. A range of scholarly works have utilized topic modeling to reveal latent thematic structures within such texts [Resch2018] [Hannigan2019]. However, the resulting topic representations are rarely contextualized with additional data modalities. Early efforts to analyze geo-social textual media in a multimodal setting typically employed sequential workflows, wherein each modality was processed in isolation through distinct modules [yin2011geographical].
While these methods have demonstrated utility in various applications, their modular and linear design often hinders the ability to capture complex interdependencies between modalities. Furthermore, this sequential structure can introduce dependencies between processing stages, rendering the overall output sensitive to the order in which analyses are performed. More recent approaches introduced advancements toward integrating multiple modalities such as time, geographic space, semantic topics, and sentiments in the analysis of social media [gonzalez2024neural] [hanny2025]. However, they still depend on numerous isolated components and fragmented intermediate processing steps, falling short of delivering a truly end-to-end processing methodology. Nonetheless, inherent heterogeneity of input modalities in this case has already added to the complexity of the task. For example, text is symbolic and sequential, while location data is numeric and geospatial, often resulting in misaligned feature representations and inconsistencies during fusion. The lack of a unified machine learning model constrains cross-modal interactions and introduces fragile interdependencies, whereby variations in preprocessing or processing order can significantly impact both performance and interpretability.
This paper presents a new methodology for multimodal geospatial machine learning, that jointly embeds semantic and geospatial information in a unified graph-based space, using end-to-end learning with a composite loss to generate clusters that are both semantically coherent and spatially interpretable. The proposed methodology is application-agnostic and can, without loss of generality, be adapted to diverse forms of multimodal data integration, making it well-suited for a wide range of data fusion scenarios and downstream tasks across domains. To capture cross-modal relationships in a holistic and structured manner, we construct a unified multimodal embedding space based on a graph structure, where we can adapt diverse data types, including text, geolocation, and optionally emotions or temporal signals, into a shared representation space. In this formulation, nodes in the graph represent user-generated post embeddings, while edges encode relational signals such as semantic similarity, geographic proximity, or user relations. This enables the model to attend to multimodal interactions in a context-aware fashion.
At the core of our methodology, we train an artificial neural network in an unsupervised manner to aggregate information from a node’s neighbors, where a modified message passing mechanism allows scaling by the corresponding edge weights. A key innovation of our model is the use of a composite loss function inspired by reference-less clustering evaluation approaches. To enable robust and discriminative representation learning, the loss integrates contrastive, coherence, and alignment objectives. The contrastive component preserves local semantic structure by pulling similar nodes closer to each other in the embedding space while pushing dissimilar ones apart. The coherence objective encourages the formation of well-defined clusters by maximizing intra-cluster similarity and inter-cluster separation. At the same time, the alignment term stabilizes training by ensuring each node remains close to the centroid of its assigned cluster, thus enhancing representational consistency. By embedding these objectives directly into the learning process, the model is able to discover clusters that are both semantically meaningful and spatially coherent. Our approach is specifically tailored to support disaster management efforts. In emergencies triggered by natural hazards such as floods, geo-social media provides rich multimodal data from affected communities, enhancing situational awareness for first responders. Through a case study of the 2017 Hurricane Harvey in the USA, we demonstrate how geographically fine-grained multimodal clusters of social media posts closely align with the disaster’s observed impacts, illustrating the practical value of our method in real-world crisis scenarios. The framework itself is designed to be application-agnostic and, without loss of generality, can be adapted to various forms of multimodal data integration. It is well-suited for diverse data fusion scenarios and downstream tasks across different domains. Empirical evaluations demonstrate that our approach outperforms existing baselines in topic coherence, spatial compactness, and interpretability, providing a scalable and principled framework for multimodal social media analysis.
In summary, this work makes the following distinct contributions:
  • Introducing a multimodal graph-based framework that seamlessly integrates multiple modalities (specifically semantic and geospatial data) into a unified representation space.
  • Proposing a domain-agnostic methodology for multimodal data analysis, underpinned by two unsupervised architectures, MonoGraph and MultiGraph, that adapt to diverse data characteristics and application needs, positioning the approach as a versatile framework for research and practice in fields such as disaster management, public health, and social media analytics.
  • Developing a composite loss function that combines contrastive, coherence, and alignment objectives, ensuring the generation of semantically coherent and spatially interpretable clusters.
  • Demonstrating the effectiveness of our approach through a real-world disaster case study, where geographically fine-grained multimodal clusters of social media posts closely align with observed impacts.
2 Related work
To contextualize our work, we first examine general topic modeling techniques and their geospatial adaptations, specifically tailored for analyzing social media data. We then delve into graph-based topic modeling approaches, emphasizing their strengths in capturing intricate relationships within the data.
2.1 Topic Modeling
Topic modeling is a widely used technique for discovering latent semantic topics within large document collections, serving as a variant of automated content analysis that assigns one or more “topic codes” to each document, thereby providing a structured thematic summary [hoyle2022content]. Among the foundational methods, Latent Dirichlet Allocation (LDA) [blei2003latent] introduced a fully generative probabilistic framework in which each document is modeled as a mixture over latent topics, and each topic is represented as a multinomial distribution over words. The generative idea of LDA has been extended to incorporate additional dimensions such as sentiment analysis [Lin2009Joint], temporal dynamics [Blei2006Dynamic], and social media user information [zhao2011comparing], enriching the interpretability and applicability of topic models across domains. However, LDA and its variants fundamentally rely on the bag-of-words assumption, which ignores word order and semantic relationships between words, limiting their effectiveness in handling short, noisy, or highly contextual texts such as tweets and social media posts [hong2010empirical]. To overcome these limitations, embedding-based topic modeling techniques have recently emerged. Dieng et al. [dieng2020topic] introduced the Embedded Topic Model (ETM), which integrates pre-trained word and sentence embeddings with LDA’s generative framework to better capture semantic relationships. Sia et al. [sia2020neural] further demonstrated that clustering high-dimensional semantic embedding vectors, derived from neural language models, can serve as a powerful alternative strategy, achieving performance comparable to traditional topic models. Building on these embedding-based approaches, Angelov [angelov2020top2vec] proposed Top2Vec, which leverages document embeddings and clustering algorithms to identify topics without requiring the number of topics to be pre-specified. This method utilizes techniques like Uniform Manifold Approximation and Projection (UMAP) and Hierarchical Density-Based Spatial Clustering of Applications with Noise (HDBSCAN) for dimensionality reduction and flexible clustering. Grootendorst [grootendorst2022bertopic] further advanced this line of work with BERTopic, which incorporates transformer-based contextual embeddings (e.g. from BERT) combined with refined clustering strategies, consistently outperforming traditional LDA and Top2Vec in both coherence and interpretability. Despite their progress, these models generally follow a sequential and fragmented framework and often struggle to balance semantic accuracy, interpretability, and scalability, particularly when applied to the diverse and noisy nature of social media data.
2.2 Spatio-semantic Topic Modeling
As location-aware social media platforms proliferate, spatially grounded topic modeling has gained increasing attention. This line of work focuses on uncovering geographically coherent semantic patterns by integrating geospatial signals into the semantic topic modeling process. Early efforts extended classical models like LDA with spatial priors. For example, Spatial LDA (sLDA) [Yuan2013Who] and Geo-aware Topic Models [Hong2012Discovering] include location-based information as latent variables or observed metadata, enabling the discovery of region-specific topics. Similar models have been applied to a variety of domains, such as understanding natural disasters [Resch2018], flood mapping [Huang2018Flood], and epidemic monitoring [Stolerman2023Disease]. However, most spatial topic models treat location and text independently or as loosely coupled variables, which often leads to weak cross-modal alignment. In addition, spatial topic models are frequently built upon modular architectures, making them difficult to scale or generalize. More recent multimodal approaches attempt to combine geolocation, text, and other modalities, often through hybrid processing pipelines [gonzalez2024neural] [hanny2025]. While these methods capture richer context, they extract each modality feature in isolation and then merge features post hoc, limiting their ability to jointly reason across modalities. Furthermore, lack of unified data space, and the representational gap between e.g. symbolic (textual) and continuous (geospatial) data introduces alignment challenges and interpretability issues in downstream clustering or topic interpretation.
2.3 Graph-based Topic Modeling
Several studies explored graph-based approaches to better capture semantic relationships and improve topic coherence. For example, Wang et al. [wang2015community] constructed word co-occurrence graphs from corpora and applied community detection algorithms, such as the Louvain method, to identify topical clusters, demonstrating superior performance on short text datasets. Similarly, Zhao et al. [zhao2018graph] proposed GSTM, a graph-based short text topic model that utilizes word co-occurrence graphs and community detection to improve topic coherence on sparse data. Beyond purely graph-based clustering, Graph Topic Models (GTM) have integrated Graph Neural Networks (GNNs) with topic modeling to jointly learn from document-word heterogeneous graphs, capturing both network structure and semantic information [zhou2020]. However, despite their ability to incorporate relational structure, GTMs still struggles to fully capture deeper or long-range semantic dependencies, limiting their effectiveness in modeling more abstract or nuanced topics. Earlier work by Newman and Girvan [newman2006finding] established a theoretical connection between community detection in bipartite document-word networks and topic model inference, highlighting the natural synergy between network science and probabilistic topic discovery. Extensions of LDA incorporating graph structures (such as Graph-LDA) have also been proposed to embed word similarity relations directly within the probabilistic framework, yielding more semantically coherent topics [doshi2015]. More recently, Graph2Topic [graph2topic2023] presented a topic modeling framework that leverags graph clustering methods on embedding-based sentence representations. This approach improves the quality and interpretability of topics without requiring a predefined number of topics, demonstrating the practical benefits of graph-embedding integration in topic discovery. Overall, graph-based topic modeling represents a promising direction that exploits the structural properties of the input data, allowing to integrate advances in network science and deep learning to overcome limitations of traditional models.
3 Methodology
Our approach leverages a set of geotagged social media posts, each represented by a semantic embedding $h_{i}$ (derived from the textual content) and a geographic coordinate $l_{i}$. To obtain rich semantic representations, we use Sentence-BERT (SBERT), which fine-tunes transformer-based language models in a Siamese architecture. This structure enables the model to directly optimize for sentence-level similarity, and has been shown in follow-up studies to effectively capture semantic proximity in short, informal texts such as tweets [Reimers2019]. To facilitate joint reasoning over these heterogeneous data types, we convert all modalities into a unified graph structure, serving as a common representation space. Graphs offer a flexible and expressive framework for capturing both local and global relationships within and across modalities. This design enables the neural model to propagate information through neighborhoods in a context-aware and modality-aware manner, something that sequential or flat representations cannot achieve effectively. While this study focuses on semantic and geographic signals, our methodology is not limited to these two modalities; without loss of generality, it can be extended to incorporate additional modalities such as temporal, sentiment, or even visual information, provided they can be represented in a graph-compatible format. We explore two complementary graph construction strategies. In the MonoGraph formulation, we define a unified graph $\mathcal{G}_{u}=(\mathcal{V},\mathcal{E}_{u})$, where edge weights integrate both modalities:
$w_{ij}=\mu\cdot\cos(\mathbf{h}_{i},\mathbf{h}_{j})+\theta\cdot\frac{1}{1+d(\mathbf{l}_{i},\mathbf{l}_{j})},$
where $h_{n}$ are the semantic embeddings and $l_{n}$ are the geographic coordinates, and $\mu$ and $\theta$ control the relative contributions of semantic and geographic similarities, respectively. This fused graph is then processed by a single Graph Convolutional Neural Network (GCN) encoder that learns joint representations capturing both semantic and spatial proximity. In the MultiGraph formulation, we construct two independent graphs over the same node set $\mathcal{V}$: a semantic graph $\mathcal{G}_{s}=(\mathcal{V},\mathcal{E}_{s})$, with edge weights:
$w_{ij}^{(s)}=\cos(\mathbf{h}_{i},\mathbf{h}_{j}),$
and a geographic graph $\mathcal{G}_{g}=(\mathcal{V},\mathcal{E}_{g})$, with edge weights:
$w_{ij}^{(g)}=\frac{1}{1+d_{\mathrm{hav}}(\mathbf{p}_{i},\mathbf{p}_{j})},$
where $\mathbf{p}_{i}=(\phi_{i},\lambda_{i})$ denotes the latitude–longitude coordinates of node $i$ in degrees under EPSG:4326, and $d_{\mathrm{hav}}$ is the great-circle (Haversine) distance given by:
$d_{\mathrm{hav}}(\mathbf{p}_{i},\mathbf{p}_{j})=2R\arcsin\left(\sqrt{\sin^{2}\left(\frac{\phi_{j}-\phi_{i}}{2}\right)+\cos(\phi_{i})\cos(\phi_{j})\sin^{2}\left(\frac{\lambda_{j}-\lambda_{i}}{2}\right)}\right),$
where $R$ is the Earth’s radius (e.g., $R\approx 6371\,\mathrm{km}$), and all angular quantities $\phi$ and $\lambda$ are expressed in radians before computation. Similarly, this graph is processed with a dedicated GCN network as well.
Figure 1: Layout of the two models: MonoGraph (a), and MultiGraph (b)
3.1 Network Architectures
A schematic overview of both architectures is provided in Fig. 1. We employ Graph Convolutional Networks (GCNs) to model local context and neighborhood interactions across semantic and geographic dimensions in both architectures. GCNs operate via a localized message-passing scheme, updating node features according to:
$\mathbf{X}^{(l+1)}=\sigma\left(\hat{A}\mathbf{X}^{(l)}\mathbf{W}^{(l)}\right),$
where $\hat{A}$ is the normalized adjacency matrix, $\mathbf{X}^{(l)}$ denotes the feature matrix at layer $l$, $\mathbf{W}^{(l)}$ is a learnable weight matrix, and $\sigma$ is a non-linear activation function such as ReLU. Our implementation uses two GCN layers with ReLU activations. In the MultiGraph setup, dropout and layer normalization are applied between layers to enhance generalization and training stability. The MonoGraph architecture processes the unified graph with a shared GCN encoder to learn joint node embeddings. Conversely, the MultiGraph architecture builds two separate GCN branches, one for semantic and one for geographic graphs, producing semantic embeddings $\mathbf{x}_{i}^{(s)}$ and geographic embeddings $\mathbf{x}_{i}^{(g)}$ for each node $i$. These embeddings are then fused through a learnable multi-head attention mechanism. In this mechanism, the geographic embeddings serve as the queries $(Q)$, while the semantic embeddings act as both the keys $(K)$ and values $(V)$, enabling the model to selectively attend to relevant semantic features conditioned on geographic context:
$\mathbf{Z}_{i}=\text{Multi-Head}\big(\mathbf{Q}=\mathbf{x}_{i}^{(g)},\mathbf{K}=\mathbf{x}_{i}^{(s)},\mathbf{V}=\mathbf{x}_{i}^{(s)}\big),\quad\mathbf{z}_{i}=\frac{1}{h}\sum_{j=1}^{h}\mathbf{Z}_{i}^{(j)},$
where the multi-head attention layer first linearly projects $(Q,K,V)$ into $h$ subspaces (attention heads) of dimension $d_{k}=\frac{\text{embed\_dim}}{h}$, then computes scaled dot-product attention for each head as
$\text{Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})=\text{softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d_{k}}}\right)\mathbf{V},$
and finally concatenates the outputs from all attention heads into a single vector, which is then linearly transformed to produce the fused embedding. Here, $h$ denotes the number of attention heads. This fusion mechanism enables adaptive balancing of semantic and spatial information, selectively modulating the influence of nodes and edges to optimize the overall training objective.
3.2 Loss Design and Training Objective
To guide the learning of node representations, we use a composite loss that combines contrastive, coherence, and alignment objectives. The contrastive loss helps preserve local semantic structure. The coherence loss promotes cluster structure. The alignment loss further stabilizes training by encouraging each node to stay close to the centroid of its assigned cluster. By jointly optimizing these losses, the model is able to learn embeddings that capture both local neighborhood structure and global cluster organization, while also maintaining alignment with cluster prototypes. Clustering is performed using the K-means algorithm with a fixed random seed and fixed cluster numbers to ensure consistency across training epochs, although any other clustering algorithm, differentiable or not, can be used without modifying the core framework.
Coherence Loss: The coherence loss consists of two components: intra-cluster similarity, which ensures topics are internally consistent, and inter-cluster dissimilarity, which discourages overlap between topics.
Intra-cluster Similarity: The intra-cluster similarity measures how similar the embeddings of nodes within the same cluster are to each other. Intuitively, nodes that belong to the same community should have similar representations, as they share common structural or semantic properties. This similarity is commonly quantified using normalized pairwise cosine similarity, which has been shown to be an effective measure for assessing the compactness of clusters, particularly for text and high-dimensional data [ghosh2006]. The cosine similarity captures the angular distance between embeddings, providing a metric that is robust to differences in magnitude while emphasizing the relative similarity of node features:
$\mu_{\text{intra}}=\frac{\sum_{k=1}^{K}\sum_{\begin{subarray}{c}i,j\in C_{k}\\ i\neq j\end{subarray}}\cos(\mathbf{z}_{i},\mathbf{z}_{j})\cdot\mathbf{1}_{\{C(i)=C(j)\}}}{\sum_{k=1}^{K}\binom{|C_{k}|}{2}}$
where $K$ is the number of clusters, and $C_{k}$ denotes the set of nodes in cluster $k$. The denominator normalizes the similarity by the number of node pairs in each cluster. Cosine similarity is computed as:
$\cos(\mathbf{z}_{i},\mathbf{z}_{j})=\frac{\mathbf{z}_{i}\cdot\mathbf{z}_{j}}{\|\mathbf{z}_{i}\|_{2}\|\mathbf{z}_{j}\|_{2}}$
Here, $\mathbf{z}_{i}$ and $\mathbf{z}_{j}$ represent the feature embeddings of nodes $i$ and $j$, respectively, and are normalized by their L2 norms to ensure that only the direction (rather than magnitude) of the vectors influences the similarity measure.
Maximizing intra-cluster similarity encourages the model to learn node embeddings that are close to one another within the same cluster. This promotes compactness and internal consistency in the embedding space, improving the semantic or spatial cohesion of discovered clusters. In this context, higher values of intra-cluster similarity, typically close to 1, indicate tight and cohesive clusters, while values near 0 or -1 suggest poorly defined clusters.
Inter-cluster Similarity: The inter-cluster similarity captures how similar node embeddings are across different clusters. Ideally, embeddings from separate clusters should be dissimilar. We define this as:
$\mu_{\text{inter}}=\frac{\sum_{\begin{subarray}{c}i,j\in P_{\text{inter}}\\ C(i)\neq C(j)\end{subarray}}\cos(\mathbf{z}_{i},\mathbf{z}_{j})\cdot\mathbf{1}_{\{C(i)\neq C(j)\}}}{\sum_{i\neq j}\mathbf{1}_{\{C(i)\neq C(j)\}}}$
Here, $P_{\text{inter}}$ denotes all node pairs from distinct clusters. High inter-cluster similarity values would indicate insufficient separation between clusters, while values close to -1 reflect distinct and well-separated clusters. So, ideally we aim for low inter-cluster similarity, with values closer to -1 indicating optimal separation between clusters.
To jointly optimize intra-cluster compactness and inter-cluster separation, we define the coherence loss:
$\mathcal{L}_{\text{coherence}}=-\left(\frac{\sum_{k=1}^{K}\sum_{\begin{subarray}{c}i,j\in C_{k}\\ i\neq j\end{subarray}}\cos(\mathbf{z}_{i},\mathbf{z}_{j})}{\sum_{k=1}^{K}\binom{|C_{k}|}{2}}-\lambda\cdot\frac{\sum_{\begin{subarray}{c}i,j\in P_{\text{inter}}\\ C(i)\neq C(j)\end{subarray}}\cos(\mathbf{z}_{i},\mathbf{z}_{j})}{\sum_{i\neq j}\mathbf{1}_{\{C(i)\neq C(j)\}}}\right)$
where $\lambda$ is a hyperparameter that controls the trade-off between maximizing intra-cluster similarity and minimizing inter-cluster similarity. This loss function combines both clustering principles into a single differentiable objective. The negative sign transforms the coherence score into a minimization problem, consistent with standard optimization procedures. Minimizing $\mathcal{L}_{\text{coherence}}$ leads to embedding spaces where nodes within the same cluster are tightly grouped, while nodes from different clusters are well-separated, enhancing both interpretability and performance of downstream clustering tasks.
Contrastive Loss: To preserve local semantic or structural neighborhoods, we use a contrastive loss that pulls embeddings of similar nodes closer and pushes dissimilar ones apart. For a batch of positive and negative node pairs, the contrastive objective is defined as:
$\mathcal{L}_{\text{contrast}}=-\sum_{(i,j)\in\mathcal{P}}\log\frac{\exp(\cos(\mathbf{z}_{i},\mathbf{z}_{j})/\tau)}{\sum_{(i,k)\in\mathcal{N}_{i}}\exp(\cos(\mathbf{z}_{i},\mathbf{z}_{k})/\tau)}$
where $\mathcal{P}$ denotes the set of positive node pairs (e.g., semantically or spatially close), $\mathcal{N}_{i}$ includes both positives and negatives for node $i$, and $\tau$ is a temperature hyperparameter controlling the sharpness of the distribution. Contrastive loss enforces local consistency in the learned embeddings, ensuring that nodes with similar contexts (semantic or spatial) are embedded close together. This regularizes the representation space and improves generalization, especially in heterogeneous graphs.
Alignment Loss: The alignment loss encourages node embeddings to remain close to the centroid of their assigned cluster. This constraint stabilizes training and reduces noise by enforcing tighter adherence to cluster prototypes. The loss is defined as:
$\mathcal{L}_{\text{align}}=\frac{1}{N}\sum_{i=1}^{N}\left\|\mathbf{z}_{i}-\mathbf{c}_{C(i)}\right\|_{2}^{2}$
where $\mathbf{z}_{i}$ is the embedding of node $i$, and $\mathbf{c}_{C(i)}$ is the centroid of cluster $C(i)$, computed as the mean embedding of all nodes in that cluster. This loss helps to maintain cluster cohesion by anchoring each node close to its corresponding cluster center. When combined with coherence and contrastive losses, alignment loss enhances the structural integrity of the latent space, especially in high-noise environments or early training stages.
Total Loss Objective: To effectively learn cluster-aware and structure-preserving embeddings, we combine all three objectives into a unified loss function. The total loss is defined as:
$\mathcal{L}_{\text{total}}=\alpha\cdot\mathcal{L}_{\text{contrast}}+\beta\cdot\mathcal{L}_{\text{coherence}}+\gamma\cdot\mathcal{L}_{\text{align}}$
where $\alpha$, $\beta$, and $\gamma$ are weighting coefficients that control the relative contribution of each component. This composite formulation allows the model to jointly optimize for local consistency, global cluster structure, and prototype alignment.
3.3 Embedded Features Clustering
After obtaining the embedded features produced by the final layer of the GCN encoder, we apply unsupervised clustering to the resulting embeddings $\mathbf{z}_{i}$. Clustering helps to reveal latent topic or community structures in the joint semantic-geographic representation. We use Spectral Clustering, which first computes a similarity matrix that reflects the pairwise relationships between data points, and then utilizes the eigenvectors of the graph Laplacian to partition the data. This method is particularly effective for high-dimensional and non-convex data distributions, such as those found in graph-based embeddings [von2007]. Likewise, we fix the number of clusters $K=10$ to stabilize the process and reduce variability, ensuring consistent topic identification across runs. This choice also enhances reproducibility by maintaining the same clustering configuration.
4 Validation
A comprehensive validation was conducted to rigorously assess the performance and generalization capability of the proposed methodology. We employed a multi-faceted experimental procedure for this purpose. First, we evaluated the models’ ability to capture coherent patterns and embed the social media data into clusters that are cohesive and distinct, corresponding to high intra-cluster similarity and low inter-cluster similarity. Next, we benchmarked their performance against the most comparable method identified in the literature, as well as a sequential baseline, across four diverse datasets. Finally, the utility of the proposed models was demonstrated through a case study on the 2017 Hurricane Harvey in the United States, in which we evaluated the interpretability and the spatial autocorrelation patterns found in the relevant topics, which are proven to be significant for disaster management applications.
4.1 Experimental Setup
Datasets: As the models were designed specifically for the analysis of social media posts in disaster situations, we validated our proposed models using four tweet datasets related to distinct real-world disaster scenarios: (1) the 2014 Napa earthquake [brocher2015], which occurred on August 24, 2014. The geographical bounds focus on the Napa Valley region in California, covering 4,287 km². (2) Hurricane Harvey in 2017 [noaa2017], a Category 4 hurricane that impacted the southeastern Texas region from August 25 to September 7, 2017. The dataset reflects a large spatial extent of 26,736 km², particularly around Houston. (3) The 2021 Ahr Valley flood [koks2021], caused by heavy rainfall, which took place in July 2021. The dataset covers the 1,256 km² area of the Ahr Valley in Germany. (4) The 2023 Chile wildfires [cordero2024], which affected central and southern Chile and were most intense in February 2023. The dataset spans a vast area of 546,905 km². All the datasets were collected using Twitter’s former v1.1 streaming and recent search API endpoints, while filtering by the bounding box Polygon (Geographic Bounds), and the posting time (Time Frame), as depicted in Table 1.
As a common practice in natural language processing (NLP), we preprocessed the raw text data (tweets) before performing further analysis. This preprocessing is done to ensure consistency, remove noise, and improve the quality of the data. Specifically, we removed URLs using a regular expression and converted the text to lowercase for keyword extraction. HTML entities were unescaped, Mentions (i.e., @references) were removed, and extra whitespace was stripped. Non-alphanumeric characters were removed, followed by the elimination of numbers. For BERT-specific cleaning, Mentions were replaced with @user, URLs were converted to http, and newline characters were removed. These steps ensure that the tweets are appropriately formatted for further analysis and model training.
Table 1: Properties of the datasets used for validation. The time frame and geographic bounds were used for filtering. In addition, we computed the number of posts within each dataset and the size of the cover area.
Dataset Posts Time Frame Geographic Bounds Cover Area
Ahr Valley floods 11,177 2021-07-01 – 2021-07-31 (50.01, 6.00, 50.75, 8.00) 1,256 km²
Hurricane Harvey 28,933 2017-08-25 – 2017-09-07 (27.58, –98.77, 30.66, –90.86) 26,736 km²
Chile wildfires 100,707 2023-02-01 – 2023-02-28 (–55.95, –76.89, –17.34, –65.22) 546,905 km²
Napa earthquake 90,907 2014-08-24 – 2014-08-25 (37.11, –123.05, 38.98, –121.04) 4,287 km²
Network Training and Parametrization: The models was trained in an unsupervised manner, without access to ground truth topic or cluster labels, using the compound loss, as detailed in the Section 3.2. Input features consisted of SBERT-generated semantic embeddings for textual content and corresponding geolocation coordinates for graph construction. These modalities were used to define semantic and geographic edges in both the MonoGraph and MultiGraph formulations. Training was performed using the Adam optimizer with an initial learning rate of 1e-3. Models were trained for 50 epochs, with early stopping based on the stability of the coherence score over a held-out validation set. In the MultiGraph setup, we applied a dropout rate of 0.3 and used two Torch GCN layers per branch. The multi-head attention mechanism used for embedding fusion consisted of 4 heads and incorporated layer normalization between fusion stages. To efficiently manage memory, avoid over-sized graphs, and enable consistent training across varying dataset sizes, we adopted a strided chunking protocol. In specific, each dataset was partitioned into multiple non-overlapping chunks using a fixed stride value. Given a stride $s$, each chunk consisted of every $s$-th entry from the original dataset, starting from a different offset. Formally, we generated $s$ subsets $\{\text{chunk}_{i}\}_{i=1}^{s}$ such that:
$\text{chunk}_{i}=\text{Input}_{\text{gdf}}[i::s]$
This approach preserved the overall diversity and temporal structure of the data while enabling the model to process each chunk independently. The protocol ensures that all data points are covered, and total recovery of the original dataset was verified before training. Each chunk was used sequentially as input to the model in a looped training cycle. Hyperparameter tuning was carried out via a random search over a validation subset of the datasets. Ultimately, we set the contrastive, coherence, and alignment loss weights to $\alpha$ = 0.8, $\beta$ = 0.2, and $\gamma$ = 0.1, respectively. These choices are based on a combination of domain knowledge, empirical tuning, and the relative importance of each loss component for the task at hand. For the contrastive loss, we set the temperature parameter $\tau$ = 0.5. All experiments were executed on an NVIDIA H100 GPU.
Table 2: Clustering evaluation results across different datasets using the two models against the input referenced data
Dataset Ahr Valley Floods Hurricane Harvey Chile Wildfires Napa Earthquake
Model Inter Intra Inter Intra Inter Intra Inter Intra
Input 0.432 0.779 0.401 0.693 0.355 0.525 0.319 0.447
MonoGraph 0.046 0.846 0.045 0.817 -0.071 0.777 -0.054 0.774
MultiGraph 0.248 0.853 0.164 0.824 -0.002 0.790 -0.042 0.791
Figure 2: Similarity matrices and their corresponding 3D clustering visualizations (after dimensionality reduction) for a subset of the Napa dataset. The first row (a, b) shows the input embeddings, while the second row (c, d) and third row (e, f) display the results for MonoGraph and MultiGraph models, respectively.
4.2 Clustering Capability Evaluation
To evaluate the models’ capacity for generating semantically coherent representations, we conducted clustering assessments on both the original SBERT input embeddings and the output embedded features produced by each model. Following the methodology outlined in Section 3.2, we computed two key metrics: intra-cluster similarity, reflecting semantic cohesion within clusters, and inter-cluster similarity, indicating the degree of separation between them. Ideally, effective clustering is characterized by high intra-cluster cohesion and low inter-cluster similarity. Table 2 summarizes the results across the four disaster-specific datasets. The input embeddings serve as a baseline and consistently exhibit moderate cohesion but relatively poor separation, suggesting that the raw SBERT representations lack strong semantic structure. Both MonoGraph and MultiGraph yield substantial improvements. Across all datasets, inter-cluster similarity drops significantly, while intra-cluster cohesion improves. MultiGraph achieves the highest cohesion values on all datasets and attains the lowest (i.e., most negative) inter-cluster similarity scores for the Chile wildfires and Napa earthquake datasets, suggesting well-isolated and internally consistent clusters. On the Ahr Valley floods and Hurricane Harvey datasets, MonoGraph shows better (lower) inter-cluster similarity values, although its intra-cluster scores are slightly lower than MultiGraph’s, yet still markedly better than the baseline. These trends are further reinforced by the visualizations in Fig. 2. The input similarity matrix (top-left) exhibit less well-defined structure, whereas the output matrices (middle-left and bottom-left) display sharper contrast and more structured groupings. Likewise, the 3D cluster plots (right) reveal a transition from scattered distributions in the input space to more compact and semantically coherent clusters after model transformation. Low inter-cluster similarity values indicate strong dissimilarity between clusters, reflecting effective cluster separation. In summary, both models, especially MultiGraph, demonstrate a strong ability to reorganize the input features space in ways that enhance clustering quality. This improvement is crucial for downstream tasks such as topic modeling, trend detection, and spatial-semantic analysis.
4.3 Benchmark and Baseline Comparison
The objective of this evaluation was to quantitatively assess the models’ performance and their capability to generate high-quality clusters across diverse disaster scenarios. For benchmarking purposes, we compared our proposed models against the multimodal pipeline presented by Hanny et al. [hanny2025], as well as the baseline sequential workflow introduced in this study. To rule out the effect of time and sentiment features in the multimodal pipeline, we set the sentiment and time weights in the pipeline to zero during the evaluation. To ensure fair comparison, we used the same exact datasets as used in [hanny2025]. For embedding generation, we utilized the English-only all-MiniLM-L12-v2 SBERT model for the 2014 Napa earthquake and 2017 Hurricane Harvey datasets. Conversely, for the 2021 Ahr Valley flood and 2023 Chile wildfires datasets, we adopted the multilingual distiluse-base-multilingual-cased-v1 model, consistent with the approach in [hanny2025]. To ensure a congruent comparison across models, we chose Topic Quality (TQ), a widely used metric in the community. TQ effectively evaluates the coherence and relevance of topics, making it a reliable measure for assessing the quality of topic models. It is computed based on the top 10 keywords per topic and defined as the product of Topic Coherence (TC) and Topic Diversity (TD). We employed a TF-IDF-based keyword extraction method: For each cluster, the corresponding subset of documents was aggregated, and TF-IDF scores were calculated over this localized corpus. The top 10 highest-scoring terms were selected as representative keywords.
To compute the TQ score, we first constructed a document-term matrix using CountVectorizer, then estimated word and word-pair frequencies across the full corpus. For each topic, pairwise Positive Pointwise Mutual Information (PPMI) between all keyword pairs was calculated and normalized by the negative log joint probability to obtain the normalized variant (NPPMI) [bouma2009] [jurafsky2023]. Formally, the normalized PPMI between two words $w_{i}$ and $w_{j}$ is given by
$\text{NPPMI}(w_{i},w_{j})=\frac{\max(0,\log\frac{P(w_{i},w_{j})}{P(w_{i})P(w_{j})})}{-\log P(w_{i},w_{j})},$
where probabilities are estimated from co-occurrence counts over the corpus. The coherence of a topic is the average NPPMI over all $\binom{10}{2}$ word pairs in the topic, and the overall Topic Coherence is the mean across all valid topics. Topic Diversity quantifies how distinct the discovered topics are by measuring the proportion of unique keywords across all topics. It is calculated as
$\text{TD}=\frac{U}{k\times|T|},$
where $U$ denotes the number of unique words in the top-$k$ keywords across all topics, $k$ is the number of keywords per topic, and $|T|$ is the total number of topics. This penalizes repeated keywords across topics and rewards diversity in word usage. Finally, Topic Quality is computed as the product of the two metrics:
$\text{TQ}=\text{TC}\times\text{TD},$
which favors models that produce both semantically coherent and lexically diverse clusters. All metrics are computed using the same preprocessing pipeline for consistency across evaluated methods.
4.4 TQ Results and Analysis
The Topic Quality (TQ) scores reported in Table 3 clearly demonstrate the effectiveness of our proposed models in generating high-quality topic clusters across diverse disaster-related datasets. The MultiGraph consistently achieves the highest TQ scores in three out of four scenarios, the Ahr Valley floods (0.415), Hurricane Harvey (0.384), and the Napa earthquake (0.195), demonstrating its robustness and adaptability to datasets with varying linguistic, geographic, and topical characteristics. This superior performance can be attributed to its cross-attention mechanism, which enables selective regulation of individual modalities and more effective capture of both global semantic patterns and local topical signals. On the Chile wildfires dataset, however, the MonoGraph slightly outperforms the MultiGraph (0.169 vs. 0.161), with both models performing comparably well overall. This suggests that while the MultiGraph generally excels, the benefits of increased model complexity may vary depending on the nature and quality of the data. In particular, contexts characterized by fewer high-quality labeled signals or more multilingual noise may favor the simpler and more direct MonoGraph architecture, which consistently outperforms baseline methods across all datasets and yields stable, focused clusters despite its relative simplicity.
Table 3: Topic quality scores (TQ) for outputs from our two models (MonoGraph and MultiGraph), the method by Hanny et al. [hanny2025], and the baseline sequential workflow across four datasets.
Dataset MonoGraph MultiGraph Hanny et al. [hanny2025] Sequential Workfollow
Ahr Valley Floods 0.383 0.415 0.235 0.029
Hurricane Harvey 0.368 0.384 0.222 0.042
Chile Wildfires 0.169 0.161 0.117 0.034
Napa Earthquake 0.191 0.195 0.143 0.030
The multimodal pipeline [hanny2025] shows limited performance (TQ ranging from 0.117 to 0.235), likely due to the removal of sentiment and temporal features during evaluation. These components, while integral to the original design, appear essential to the pipeline’s effectiveness. While the model presented in this paper was originally designed to incorporate sentiment and temporal features, these were deliberately neutralized in our evaluation to isolate textual topic modeling performance. The resulting drop in quality highlights the fact that the pipeline’s heavy reliance on auxiliary modalities may limit its effectiveness in settings where such features are unavailable, noisy, or intentionally excluded. The sequential workflow performs the worst overall, with TQ scores falling below 0.05 across all datasets. This approach suffers from treating the stages of the modeling pipeline, embedding, clustering, and keyword extraction, as independent components, which results in fragmented topic structure and poor alignment between cluster semantics and extracted keywords. The absence of feedback or interaction between stages appears to hinder its ability to produce globally meaningful topic representations.
Figure 3: Topic-word distributions for a subset of the Hurricane Harvey dataset, applying MonoGraph (top) and the MultiGraph (bottom) models.
5 Case Study Evaluation
A key motivation behind the design of our models was to ensure interpretability for real-world applications. To complement the experiments in Section 4, we conducted a case study to showcase their utility in a disaster management context. We used a subset of Hurricane Harvey dataset, a tropical storm that occurred in 2017 and hit the state of Texas in the USA on August 25 [noaa2017]. The dataset comprises 28,933 geo-located tweets posted within and around the area of Texas, obtained by filtering tweets based on Geographic Bounds defined by the polygon $\text{Polygon}(-98.77,27.58;-90.86,27.58;-90.86,30.66;-98.77,30.66;-98.77,27.58)$ and the Time Frame from August 25 to September 7, 2017.
5.1 Topic Interpretation
To gain qualitative insights into the latent structure of the discovered clusters, we analyzed both the semantic coherence and spatial distribution of topics generated by the MonoGraph and MultiGraph models applied to a subset of the dataset. The topic-word distributions, summarized in Fig. 3, reveal semantically coherent and lexically diverse clusters that reflect key aspects of disaster-related discourse, and several topics capture geographically grounded and event-focused narratives. In particular, Topic 0 groups terms such as “houston,” “texas,” “harvey,” “hurricaneharvey,” “hurricane” and “flood,” reflecting a strong spatial anchoring of disaster-related conversations. As reflected in the coresponding spatial visualization in Fig. 4, the distribution of this topic closely follows the hurricane’s path, with dense tweet concentrations in Houston and along the Texas coastline. This alignment suggests that the models successfully identify localized, event-driven public responses directly associated with Hurricane Harvey’s impact zone. Another distinct cluster is Topic 2 in MonoGraph and Topic 1 and in MultiGraph, centered on traffic-related concerns, including terms like “traffic,” “accident,” “mins,” “delay,” and “stop.” the spatial maps reveal concentrated tweet activity along major transport corridors in southeastern Texas, likely corresponding to road closures, congestion, and transportation disruptions resulting from flooding and evacuation efforts.
Figure 4: Spatial distribution (KDE) of the Topic-words for a subset of the Hurricane Harvey dataset, applying MonoGraph (top) and the MultiGraph (bottom) models.
Other topics represent broader or more indirect engagement with the event. Topic 8 in MonoGraph and Topic 5 and in MultiGraph, captured meteorological observations, with keywords such as “temperature,” “humidity,” “clouds,” “rain,” and “wind,” are more evenly distributed across central Texas, including Austin and San Antonio. These spatial patterns suggest contributions from users outside the most severely affected areas, often in the form of commentary or weather updates rather than first-hand impact reports. Weather-related discourse is further reflected in Topics 2, 7, 5 and 9, in MonoGraph and Topics 3, 4, 9, 6 and 8 in MultiGraph, featuring overlapping terms such as “pressure,” “gust,” “humidity,” “wind,” and “clouds.” While this lexical overlap reduces topic distinctiveness, it underscores the recurring presence of weather-related signals across the corpus, a common characteristic in disaster-related social media data. This also illustrates the trade-off of fixing the number of topics ($K=10$), which facilitates stability but can lead to partial redundancy. Finally, content-sharing and localized observational behavior emerge as a consistent pattern. For example, certain clusters contain terms such as “posted,” “photo,” and place names like “meyerland,” “katy,” and “downtown,” suggesting the dissemination of real-time visual documentation and location-tagged updates by affected residents. This pattern reflects well-known crisis communication behavior on social media, where users share on-the-ground images to inform others and raise awareness.
Figure 5: Moran’s I plots for MonoGraph (top) and MultiGraph (bottom) Models.
5.2 Global Autocorrelation Interpretation
To explore the global distribution of tweet activity associated with different topics, we applied Moran’s I analysis. This method is particularly beneficial as it helps to identify spatial autocorrelation, revealing whether tweet activity related to specific topics is clustered in certain geographic regions or distributed randomly around. The metric ranges from -1 (perfect dispersion) to +1 (perfect clustering), with 0 indicating randomness. A p-value << 0.05 suggests the observed spatial pattern is statistically significant and unlikely due to chance. The related scatter plots and their associated statistics, demonstrated in Fig. 5 complement the spatial intensity patterns visualized in Fig. 4. In the MultiGraph model, Topic 0 shows a Moran’s I of 0.076 (p = 0.001), suggesting weak but statistically significant clustering. Tweets on this topic tend to occur more frequently near other tweets discussing the same theme than by chance, although the clustering is not very strong. The slight clustering might indicate localized interest or discussions related to this topic in specific regions. Topic 2 has a Moran’s I of 0.178 (p = 0.001), indicating weak but significant spatial clustering. The positive value suggests a tendency for tweets on this topic to be more closely located in certain regions. This could reflect a localized trend where users in particular areas are more engaged with or affected by the theme of the topic. Topic 5 has a Moran’s I of -0.032 (p = 0.001), a slightly negative Moran’s I indicating weak spatial dispersion, with tweets being more likely to occur near unrelated content. This suggests a dispersed pattern of topic-related tweets, though the effect is weak. The scattered distribution might indicate that this topic is more widespread, with no strong geographic concentration.
In the MonoGraph model, Topic 0 shows a Moran’s I of 0.065 (p = 0.001), which is similar to the MultiGraph model, showing weak clustering, though the clustering effect is slightly weaker than in the MultiGraph, with tweets on this topic being somewhat more scattered. This could imply a broader discussion, not concentrated in specific areas, though still more frequent in certain locations compared to random. Topic 1 has a Moran’s I of 0.201 (p = 0.001), indicating moderate clustering. Tweets on this topic show a clear tendency to be geographically concentrated in certain regions, which suggests a more localized theme. The higher Moran’s I value suggests that the topic is more likely to spark concentrated discussions within specific areas, potentially reflecting regional events or discussions. Topic 8 has a Moran’s I of -0.035 (p = 0.001), indicating a weakly dispersed distribution of tweets, showing no significant clustering and pointing to random spatial dispersion. The lack of clustering suggests that the topic has less geographic relevance or is more universally spread, with little concentration in specific regions.
Overall, most topics exhibit weak spatial clustering, as indicated by positive but modest Moran’s I values for topics 0, 1, and 2, reflecting slight regional concentration without distinct spatial boundaries. Topic 0, directly related to disasters, shows some geographic focus, while topics 1 and 2, indirectly related to disaster events, display weaker clustering. In the MonoGraph model, Topic 1 demonstrates stronger clustering (Moran’s I = 0.201), suggesting a more localized concentration. Topics 5 and 8 show weak spatial dispersion, with low or negative Moran’s I values, indicating a more even distribution across geographic regions and a lack of clear spatial patterns.
Figure 6: Spatial autocorrelation analysis (the Gi* maps (top) and LISA maps (bottom)) for the MonoGraph model.
Figure 7: Spatial autocorrelation analysis (the Gi* maps (top) and LISA maps (bottom)) for the MultiGraph model.
5.3 Spatial Autocorrelation Analysis
To further investigate the spatial (local) dynamics of identified topics, we applied the Getis-Ord Gi* statistic and Local Indicators of Spatial Association (LISA) to both the MonoGraph outputs. From a disaster management perspective, combining Gi* with LISA provides a powerful way to interpret social media–derived topics in both their intensity and spatial context. Gi* hotspots highlight areas of concentrated disaster-related activity, helping to quickly pinpoint zones of immediate impact for targeted relief. LISA adds a layer of context by classifying regions into high–high, low–low, high–low, and low–high clusters, showing not only where activity is high but also how it relates to surrounding areas. This combined approach enables planners to identify both core impact zones and peripheral but relevant areas, anticipate spillover effects, and spot isolated needs. In our results, the MultiGraph (Fig. 7) tends to produce more localized clusters, which may be better suited for tactical, on-the-ground coordination, while the MonoGraph model’s broader patterns (Fig. 6) can inform strategic, statewide planning. For Topic 0 in both models, representing core Hurricane Harvey discourse, the Gi* analysis revealed statistically significant hotspots spanning Houston, the Texas Gulf Coast, and inland counties along the storm’s path, with corresponding coldspots in northern and western Texas. LISA clustering confirmed persistent high–high (HH) clusters in these impact zones, indicating strong spatial concentration of disaster-related conversation. Such patterns can guide response agencies toward areas most actively engaged in reporting severe impacts, closely aligning with ground-truth disaster footprints.
As it can be seen in Figs. 6 and 7, traffic-related topics (MonoGraph Topic 2, MultiGraph Topic 1) produced more localized hotspots in and around metropolitan Houston and along major transport corridors. Gi* maps highlight narrowly bounded high-value clusters, while the LISA maps show HH concentrations surrounded by low–low (LL) areas, indicating geographically confined yet intense transportation disruptions. This suggests that mobility restoration efforts could be prioritized for specific corridors rather than across broad regions. Weather observation topics (MonoGraph Topic 8 MultiGraph Topic 5) exhibited scattered hotspots interspersed with wide coldspot zones. LISA classifications revealed fewer HH clusters but a notable presence of low–high (LH) outliers, counties with high weather-related discourse surrounded by low-activity neighbors. These may represent isolated monitoring stations, amateur weather observers, or unaffected communities contributing updates, which can complement official meteorological networks during disaster response.
6 Discussion and Limitations
Across all datasets, both MonoGraph and MultiGraph consistently outperformed baseline methods, achieving higher intra-cluster cohesion and lower inter-cluster similarity. Their relative strengths varied with event characteristics, suggesting sensitivity to spatial scale, discourse patterns, and data quality. Fixing the number of clusters at $K=10$ facilitated direct cross-model comparisons and system stability, but also introduced constraints. In some cases, semantically similar clusters emerged, particularly in weather-related topics, where recurring terms such as “pressure,” “humidity,” “wind” were dispersed across multiple clusters. This effect may be amplified in datasets containing unfiltered automated “weather bot” content. Empty or low-content clusters occasionally appeared due to sparse coverage, noise, or embedding collapse. The spatial distribution of events significantly shaped clustering performance. MultiGraph produced the lowest inter-cluster similarity for geographically concentrated events (Chile wildfires, Napa earthquake), benefiting from modality-specific graph encoding and cross-attention fusion. MonoGraph showed better separation in smaller or noisier AOIs (Ahr Valley floods, Hurricane Harvey), where its unified representation was more robust to diffuse or overlapping discourse. Spatially, MultiGraph generated compact, well-defined hotspots with fewer outliers, aligning clusters closely with localized disaster impacts, an asset for targeted response planning. MonoGraph captured broader spatial patterns, useful for representing large-scale narratives but with a tendency to overestimate geographic extent.
Several limitations merit consideration. First, the framework’s performance is sensitive to the quality and balance of input modalities. Second, the current implementation does not explicitly model temporal dynamics, which constrains the ability to capture and analyze the evolution of topics over time. Third, the use of a fixed number of clusters ($K$) may not reflect the true topical complexity of diverse events, suggesting the need for adaptive or non-parametric clustering strategies. Finally, MultiGraph’s scalability to real-time, high-volume streams is untested. These constraints may hinder applicability to scenarios involving extremely high-frequency data or expansive geographic coverage unless complemented by distributed computing or streaming graph processing frameworks. Despite these constraints, both architectures effectively detect geographically specific and semantically coherent topics that align with real-world hazard footprints, offering practical value for enhancing situational awareness and supporting decision-making in disaster management.
7 Conclusion
This work presented an unsupervised multimodal graph-based methodology for the semantic and spatial analysis of geo-social media data. By jointly modeling semantic embeddings and geographic coordinates through graph-based architectures, the approach addresses limitations of prior models that treat these modalities in isolation. Two architectural variants, MonoGraph and MultiGraph, were introduced to accommodate differing data characteristics and application needs. A composite loss function integrating contrastive, coherence, and alignment terms was key to learning robust, cluster-aware representations that preserve both local and global structure. Extensive evaluation on four real-world disaster datasets demonstrated consistent improvements over state-of-the-art baselines in topic coherence, spatial compactness, and overall topic quality. The MultiGraph variant, employing modality-specific encoding with attention-based fusion, achieved the strongest performance in most settings, while qualitative analysis of Hurricane Harvey (case study) confirmed the interpretability and geographic relevance of the discovered topics. The proposed methodology offers a scalable, principled approach to integrate symbolic and spatial signals for geo-social media analysis. Although the current implementation focuses on text and location, it can be extended to additional modalities such as sentiment, temporal dynamics, or imagery, enabling richer, more context-aware analyses in domains including disaster management, urban planning, and public health surveillance. Future directions include incorporating time-aware modeling, improving scalability for real-time data streams, and leveraging social network structures to enhance community-level insights. By unifying multimodal signals in an end-to-end learning model, this work advances toward more holistic and interpretable models for understanding the complex information landscapes of social media.